Circular No. 6070

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET MACHHOLZ 2 (1994o) Independent reports of a third diffuse object moving with the same

motion as comet 1994o have been received from Petr Pravec (Ondřejov, 0.65-m reflector) and Wayne Johnson (Anza, CA, 0.56-m reflector) on CCD images obtained Sept. 2.11 and 3.51 UT, respectively. Pravec noted a 1′.0 coma with little condensation, the third companion being located ∼ 43′′ southwest of the second object reported by Jäger (IAUC 6066); Johnson noted this third component as being ∼ 1 mag fainter than the second.

COMET NAKAMURA-NISHIMURA-MACHHOLZ (1994m) Improved orbital elements from MPC 23848:

T = 1994 July 12.9393 TT  ω = 123.^{o}0713
Ω = 158.9124 }2000.0
q = 1.140138 AU  i = 94.3780
1994TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Sept. 5 21^{h}56.^{m}53 + 1^{o}23.^{′}0 0.428 1.424 163.^{o}3 11.^{o}8  7.7
10 21 36.29 −13 09.4 0.495 1.470 154.9 16.9  8.1
15 21 21.18 −23 30.7 0.591 1.519 142.7 23.7  8.7
20 21 10.25 −30 34.9 0.705 1.569 132.4 28.2  9.2
25 21 02.64 −35 26.3 0.830 1.620 124.0 30.9  9.7
30 20 57.70 −38 50.8 0.961 1.673 116.9 32.3 10.1
Oct. 5 20 54.90 −41 17.4 1.095 1.726 110.8 32.8 10.6
10 20 53.84 −43 04.6 1.232 1.781 105.4 32.7 11.0
15 20 54.21 −44 24.1 1.370 1.836 100.5 32.3 11.3
20 20 55.75 −45 23.6 1.508 1.891  95.9 31.6 11.7
25 20 58.26 −46 08.3 1.645 1.947  91.7 30.7 12.0

Further total visual magnitude estimates: Aug. 27.21 UT, 7.5 (C. E. Spratt, Victoria, BC, 0.20-m reflector); 28.88, 8.3 (H. Mikuž, Cres, Slovenia, 7×50 binoculars); 30.08, 8.3 (J. E. Bortle, Stormville, NY, 20×80 binoculars); 31.21, 7.8 (Spratt, 0.10-m refractor); 31.89, 8.7 (Mikuž); Sept. 1.89, 8.8 (A. Diepvens, Balen, Belgium, 0.15-m refractor).

1994 September 3 Daniel W. E. Green